OpenStudy (anonymous):

What is the approximate distance from (2, 5) to (-2, 1)?

OpenStudy (amistre64):

subtract the points; square them, add them; then sqrt them

OpenStudy (anonymous):

\[\sqrt{(2+2)^2+(5-1)^2}\]

OpenStudy (amistre64):

(2, 5) -(-2, 1) -------- (4, 4)^2 = (16+16) = 32 sqrt(32) = sqrt(16*2) = 4 sqrt(2)

OpenStudy (amistre64):

sqrt(1), sqrt(2), sqrt(3), sqrt(4) 1 ** 2 sqrt(2) is closer to 1 so id say its a little bigger than 4

OpenStudy (anonymous):

So 5.65?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

use the distance formula!\[D=\sqrt{(X _{1}-X _{2})+(Y _{1}-Y _{2}})\]

OpenStudy (amistre64):

the distance formula still has to be stepped thru; which is why i just go to the steps :)